# Moving toward best practice when using propensity score weighting in survey observational studies

### v1

@@ Abstract
Propensity score weighting is a common method for estimating treatment effects with survey data. The method is applied to minimize confounding using measured covariates that are often different between individuals in treatment and control. However, existing literature does not reach a consensus on the optimal use of survey weights for population-level inference in the propensity score weighting analysis. Under the balancing weights framework, we provided a unified solution for incorporating survey weights in both the propensity score of estimation and the outcome regression model. We derived estimators for different target populations, including the combined, treated, controlled, and overlap populations. We provide a unified expression of the sandwich variance estimator and demonstrate that the survey-weighted estimator is asymptotically normal, as established through the theory of M-estimators. Through an extensive series of simulation studies, we examined the performance of our derived estimators and compared the results to those of alternative methods. We further carried out two case studies to illustrate the application of the different methods of propensity score analysis with complex survey data. We concluded with a discussion of our findings and provided practical guidelines for propensity score weighting analysis of observational data from complex surveys.
Keywords: complex survey, propensity score weighting, survey weights, overlap weights, covariate balance, augmented weighting estimators
@@ 1 Introduction
The use of observational data to estimate the causal effects of exposures or interventions is prevalent in medical research, particularly when randomized trials are unavailable or unethical. In many cases, researchers resort to observational data garnered through surveys with complex sampling techniques. These techniques are designed to ensure data from a sample can accurately represent the target population. In extensive surveys, these techniques may encompass complex features such as multistage sampling and deliberate oversampling of less representative subgroups.[1] The survey weights computed based on the design are integral to recovering the population features so that the differentiated selection probabilities and non-responses are appropriately accounted for. There is consensus on the importance of incorporating the survey design into the analytical process to derive causal effect estimates that are applicable to the target population.[2, 3, 4]
@@
Survey weights also play a crucial role in ensuring the accuracy and representativeness of the propensity score method, which is among the most commonly used tool for causal inference using observational data.[5] In the absence of randomization, systematic differences in baseline characteristics usually exist between exposed and unexposed individuals, leading to the issue of confounding. The propensity score methods rest on the treatment assignment process with observed covariates,[6] and with estimated propensity scores, differences in the observed baseline features across groups can be reduced so that the estimated treatment effect will no longer be attributable to pre-existing group differences. In practice, implementing propensity score methods involve a two-step procedure, including propensity score estimation and the subsequent treatment effect estimation. The latter step leverages the estimated propensity score through stratification,[7] matching,[8] or weighting.[9]
@@
The appropriate use of survey weights in stratification to obtain population causal effect estimates has been previously addressed, with a recommendation that survey weights should be used when estimating treatment effects within strata but not in the estimation of the propensity scores. [10, 2] DuGoff et al. later suggested that survey weights can carry unique design features that are not captured by observed covariates, and therefore should be included as an additional covariate in the propensity score model to enhance covariate balance and address potential bias due to unmeasured covariates.[3] Similar recommendations have also been extended to propensity score matching and weighting, where the estimation of propensity score can incorporate survey weights as an additional covariate; but in those contexts, whether to perform survey-weighted analysis in the outcome model depends on the goal of the analysis (i.e., whether a sample or the underlying population defines the target of inference).[3]. Austin et al.[11] illustrated through simulations that including survey weight as a covariate in the propensity score model showed no clear advantage over a survey-weighted propensity score model. Neither approach can consistently outperform the other regarding covariate balance and bias in the treatment effect estimation across various scenarios. Dong et al. [12] evaluated multiple strategies for incorporating survey weights into both propensity score and outcome models, recommending their use at both stages, although they observed that estimates from different strategies did not yield apparent differences in their data application. Lenis et al.[13] found through simulations that incorporating survey weights in the propensity score model does not necessarily improve the estimation of population-level treatment effect using matching, especially when sufficient population-level balance in confounders can already be achieved. They also showed through simulations that incorporating survey weights in propensity score estimation can reduce bias when the missingness in outcome is correlated with survey weights.[13]
@@
Regarding propensity score weighting, Ridgeway et al.[4] provided a theoretical justification for the use of a survey-weighted propensity score model in obtaining a consistent estimator of the population-level average treatment effect. They proved that using survey weights in both stages leads to consistent and robust population average treatment effect estimators under various generating mechanisms (e.g., survey weights are either informative or non-informative for the treatment assignment mechanism). Salerno et al. [14] recently explored a complex survey mechanism in health disparities research, where survey weights depend on the group variable of interest under comparison, such as race. They further revealed that whether the propensity score should be survey-weighted corresponds to whether the available survey weights are tied to comparison variables such as treatment or exposure. Whereas Ridgeway et al.[4] derived the results and performed simulations for the continuous outcome scenario, Yang et al.[15] extended the results to a binary outcome and evaluated the performance of different ways to incorporate survey weights in the propensity score model and the outcome model. Both Ridgeway et al.[4] and Yang et al.[15] considered the potentially differentiated practice of incorporating survey weights when estimating the average treatment effect (ATE) and the average treatment effect of the treated (ATT).
@@
In general, the emerging message from the existing literature is that the use of survey weights at the propensity score model stage and the outcome model stage can potentially be beneficial in reducing the bias of population-level inference. However, a consensus on the best practice and the underlying justifications for using survey weights to target different causal estimands of interest have not been fully elucidated. This paper thus aims to contribute to the literature on propensity score weighting in the following several ways. First, the target estimands in the prior development on survey observational data methods were limited to ATE and ATT, but not to other estimands (e.g., ATO, the average treatment effect among the overlap population[9]). We address this by supplying a unified framework that incorporates survey weights into the general framework of balancing weights.[9] Second, as a further extension, we detail the use of survey weights for augmented weighting estimators under the balancing weights framework, which can potentially improve efficiency over weighting alone by leveraging information from a posited outcome regression function. Third, we developed closed-form variance estimators that accommodate survey weights in each estimator we considered for computationally efficient inference. Lastly, given that no consensus has been reached on the best use of survey weights in propensity score weighting analysis under various real-world scenarios (such as varying levels of covariate overlap and model misspecification), we report on an extension simulation study that shed light on the relative comparisons between methods and provide recommendations to help practitioners enhance the reliability of causal evidence obtained from survey observational data.
@@
The rest of this paper is organized as follows. Section 2 defines the notation and assumptions for the potential outcomes framework under the context of survey observational data. Under the balancing weight framework, this section then derives the weighting estimator and emphasizes the role of survey weights in additional to balancing weights. We provide a unified expression of the sandwich variance estimator and demonstrate that the survey-weighted estimator is asymptotically normal, as established through the theory of M-estimators. In Section 3, we present three augmented estimators incorporating survey weights and further provide the the associated sandwich variance structure for large-sample inference. In Section 4, we conduct a full series of simulations to examine the performance of different estimators incorporating propensity score weights for different estimands with data generated from multistage clustering sampling. Section 5 compares the performance of the different estimation procedures in an application using the Early Childhood Longitudinal Study, Kindergarten Class of 1998-1999 (ESCL-K). [16] In Section 6, we present our main conclusions and provide a discussion.
@@ 2.1 Definitions and assumptions
In this section, we establish our study framework starting with the setup of finite sample and population under the context of survey observational data. Consider a finite sample of size $n$ selected from a superpopulation of size $N$. Each individual in the sample is indexed by $i$, where $i=1,\dots,n$. Assuming a $p$-dimensional vector for baseline covariates for each individual $\mathbf{X}_{i}$ that includes observed pretreatment variables, and a binary exposure variable $Z_{i}=z$ that indicates treatment ($z=1$) or control ($z=0$) group affiliation. Under the Stable Unit Treatment Value Assumption (SUTVA), the potential outcomes $Y_{i}(z)$ for $z\in\{0,1\}$ are well-defined with no interference between population units where the observed outcome for each individual in the sample is $Y_{i}=Z_{i}Y_{i}(1)+(1-Z_{i})Y_{i}(0)$. For estimating population effects using our sample, we define the survey indicator $S_{i}$ where $S_{i}=1$ denotes selected to the sample and $S_{i}=0$ denotes unselected to the sample. We assume selection positivity such that every individual in the superpopulation has a non-zero probability of selection, $10<p_{i}=p(\mathbf{x}_{i})=P(S_{i}=1\mid\mathbf{X}_{i}=\mathbf{x}_{i})<$ 1.
We assume conditional independence (or unconfoundedness) for the treatment assignment, $\{Y(0),Y(1)\}\perp Z\mid\mathbf{X}$, and for the sampling process after the exposure has occurred, $\{Y(0),Y(1)\}\perp S\mid Z,\mathbf{X}$, which implies that treatment assignment bias arises solely from $\mathbf{X}$, and selection bias from the afterwards sampling process can only be caused by $Z$ and $\mathbf{X}$. These assumptions can be relaxed to weak treatment assignment exchangeability, expressed as $E\left[Y(z)\mid\mathbf{X}\right]=E\left[Y(z)\mid Z=z,\mathbf{X}\right]$, and weak selection exchangeability, formulated as $E\left[Y(z)\mid Z=z,\mathbf{X}\right]=E\left[Y(z)\mid Z=z,\mathbf{X},S=1\right]$. These relaxed assumptions still enable the identification of causal effects and enable the generalization from the finite sample to the superpopulation, as previously noted by Salerno et al.[14] One causal estimand of interest is the average treatment effect (ATE), denoted as $\tau(\mathbf{x})=E[Y(1)-Y(0)\mid\mathbf{X}=\mathbf{x}]$. Imbens[17] differentiates population-level ATE (PATE) $\tau_{\text{PATE}}=E[Y(1)-Y(0)]$ from the sample average treatment effect (SATE), which we denote as $\tau_{\text{SATE}}=E[Y(1)-Y(0)\mid S=1]$. The population and sample level propensity scores for individual $i$ are denoted as $e_{\text{sp}}(\mathbf{x}_{i})=P(Z_{i}=1\mid\mathbf{X}_{i}=\mathbf{x}_{i})$ and $e_{\text{fp}}(\mathbf{x}_{i})=P(Z_{i}=1\mid\mathbf{X}_{i}=\mathbf{x}_{i},S_{i}=1)$ respectively. They will be equal only when sampling and treatment assignment are independent conditional on baseline covariates, i.e., $Z\perp S\mid\mathbf{X}$.
In a prospective study setup where treatment assignment occurred after the sampling, we can define the sampling probability for individual $i$ as the probability of inclusion in the study based on observed covariates, $p_{i}=P(S_{i}=1\textbar\mathbf{X}_{i}=\mathbf{x}_{i})$. In contrast, for a retrospective study where sampling occurs after exposure or treatment assignment, the sampling probability can depend on both observed covariates and treatment and is denoted as $p_{z,i}=p_{z}(\mathbf{x}_{i})=P(S_{i}=1\mid Z_{i}=z,\mathbf{X}_{i}=\mathbf{x}_{i})$. To provide further clarity, we define $p_{1}(\mathbf{x}_{i})=P(S_{i}=1\mid Z_{i}=1,\mathbf{X}_{i}=\mathbf{x}_{i})$ and $p_{0}(\mathbf{x}_{i})=P(S_{i}=1\mid Z_{i}=0,\mathbf{X}_{i}=\mathbf{x}_{i})$ as the specific sampling probabilities for the selected individuals in the treatment and control groups, respectively. When we assume conditional independence between sampling and treatment assignment, represented as $Z\perp S\mid\mathbf{X}$, the sampling probabilities for both prospective and retrospective designs align conceptually with the propensity score framework. In this framework, sampling is viewed as a separate assignment mechanism that can be adjusted for in the analysis to address potential biases. Moreover, these different sampling probabilities can then be connected through a ratio, $p_{i}=r_{z,i}p_{z,i}$, where $r_{z,i}=r_{z}(\mathbf{x}_{i})$ captures the discrepancy between propensity scores in the sample and population levels. For a treated unit ($z=1$), $r_{1}(\mathbf{x}_{i})$ equals $e_{\text{sp}}(\mathbf{x}_{i})/e_{\text{fp}}(\mathbf{x}_{i})$, the ratio of propensity score at population and sample level. For a control unit ($z=0$), $r_{0}(\mathbf{x}_{i})$ equals $(1-e_{\text{sp}}(\mathbf{x}_{i}))/(1-e_{\text{fp}}(\mathbf{x}_{i}))$. Furthermore, $r_{z}(\mathbf{x}_{i})p_{z}(\mathbf{x}_{i})$ can be used to approximate $p_{i}=P(S_{i}=1\mid\mathbf{X}_{i}=\mathbf{x}_{i})$ even when the marginalized survey weights $p_{i}$ are not directly accessible. As Salerno et al.[14] highlighted, the distinction between prospective and retrospective studies underscore two different ways of factorizing $P(S_{i}=1,Z_{i}=z\mid\mathbf{X}_{i}=\mathbf{x}_{i})$: one conditional on the covariates alone (for prospective studies) and the other conditional on both treatment and covariates (for retrospective studies). To ensure the validity of causal estimates, assumptions for positive probabilities of sampling and treatment assignment must hold at both the population and sample levels, with unconfoundedness providing the necessary ignorable conditions for both mechanisms. In this paper, we primarily focus on the retrospective study setup, which commonly occurs in survey observational data; in other words, we address the scenario where the exposure is already implemented in the target population.
Regarding the setup of propensity score weighting, we follow the framework of balancing weights [9] that characterizes a general class of weighted distributions of covariates across two treatment groups that can achieve balance. Balancing weights typically leverages the propensity score to calibrate covariates of each group towards a target population with an estimand of interest. The density of the target population is represented as $g(\mathbf{X})=f(\mathbf{X})h(\mathbf{X})$, where $f(\mathbf{X})$ signifies the marginal density of covariates within the finite population, and $h(\mathbf{X})$, known as the tilting function, characterizes the target population and estimand. The balancing weights framework articulates the estimand as a weighted average of individual-level conditional treatment effects: [9, 18]
@@
In this equation, the expectations in both the numerator and the denominator are taken over the target population. The tilting function $h(\mathbf{X})$ is a non-negative function of covariates and satisfies $E\left[h(\mathbf{X})\right]>0$.
2.2 Incorporating survey weights into propensity score weighting
Under the balancing weights framework, the tilting function $h(\mathbf{X})=g(\mathbf{X})/f(\mathbf{X})$ serves as a bridge to tilt the covariate distribution in the population $f(\mathbf{X})$ to that in the target population $g(\mathbf{X})$. Since our goal is to derive the expressions of various estimands for target populations in the superpopulation, we first derive a general form of this class of estimands $\tau_{h}(\mathbf{\mathbf{X}})$ considering the average treatment effect over the target population in the superpopulation. Under weak exchangeability in both the treatment assignment and sampling processes, Equation (1) can be expressed as,
$\tau_{h}$ $=\frac{E\left[\displaystyle\frac{h(\mathbf{X})}{p_{1}(\mathbf{X})e_{\text{sp}}(\mathbf{X})}ZY\mid S=1\right]}{\displaystyle E\left[\frac{h(\mathbf{X})}{p_{1}(\mathbf{X})e_{\text{sp}}(\mathbf{X})}Z\mid S=1\right]}-\frac{E\left[\displaystyle\frac{h(\mathbf{X})}{p_{0}(\mathbf{X})\left(1-e_{\text{sp}}(\mathbf{X})\right)}(1-Z)Y\mid S=1\right]}{E\left[\displaystyle\frac{h(\mathbf{X})}{p_{0}(\mathbf{X})\left(1-e_{\text{sp}}(\mathbf{X})\right)}(1-Z)\mid S=1\right]}.$ (2)
When the expectations are replaced with sample means, a unbiased estimator of $\tau_{h}$ in the form of a weighted difference in the outcomes between treatment groups in the target population can be expressed as

### v2

@@ Abstract
Propensity score weighting is a common method for estimating treatment effects with observational data, by addressing confounding due to measured baseline covariates. However, when the observational data sample is drawn based on a survey, the existing literature does not reach a consensus on the optimal use of survey weights for population-level causal inference. Under the balancing weights framework, we provide a unified solution for incorporating survey weights and derive a set of weighting and augmented weighting estimators for different target populations, including the combined, treated, controlled, and overlap populations. We also develop closed-form sandwich variance estimators for each estimator via the theory of M-estimators. Through an extensive series of simulation studies, we examined the performance of our estimators and compared the results to those of alternative methods. We carried out two case studies to illustrate the application of the different propensity score methods with complex survey data. We concluded with a discussion of our findings and provided some practical recommendations for propensity score weighting analysis of survey observational data.
Keywords: causal inference, propensity score weighting, survey weights, overlap weights, covariate balance, augmented weighting estimators
@@ 1 Introduction
The use of observational data to estimate the causal effects is prevalent in medical research, particularly when randomized trials are unavailable. In many cases, researchers resort to observational data garnered through surveys with complex sampling techniques. These techniques are designed to ensure data from a sample can accurately represent the external target population, and include features such as multistage sampling and deliberate oversampling of less representative subgroups.[1] The survey weights computed based on the design are integral to recovering the population features so that the differentiated selection probabilities and non-responses are appropriately accounted for. There is consensus on the importance of incorporating the survey design into the analytical process to derive causal effect estimates that are applicable to the target population.[2, 3, 4]
@@
Survey weights also play a crucial role in ensuring the accuracy and representativeness of the propensity score method, which is among the most commonly used tool for causal inference using observational data.[5] In the absence of randomization, systematic differences in baseline characteristics usually exist between exposed and unexposed individuals, leading to the issue of confounding. The propensity score methods act on the treatment assignment mechanism with measured baseline covariates,[6] and with estimated propensity scores, differences in the observed confounders across groups can be reduced so that the estimated treatment effect will no longer be attributable to pre-existing group differences. In practice, implementing propensity score methods involve a two-step procedure, including propensity score estimation and the subsequent treatment effect estimation. The latter step leverages the estimated propensity score through stratification,[7] matching,[8] or weighting.[9]
@@
The appropriate use of survey weights in stratification to obtain population causal effect estimates has been previously addressed, with a recommendation that survey weights should be used when estimating treatment effects within strata but not in the estimation of the propensity scores. [10, 2] DuGoff et al.[3] later suggested that survey weights can carry unique design features that are not captured by observed covariates, and therefore should be included as an additional covariate in the propensity score model to enhance covariate balance and address potential bias due to unmeasured covariates. Similar recommendations have also been extended to propensity score matching and weighting, where the estimation of propensity score can incorporate survey weights as an additional covariate; but in those contexts, whether to perform survey-weighted analysis in the outcome model depends on the goal of the analysis (i.e., whether a sample or the underlying population defines the target of inference).[3] Austin et al.[11] illustrated through simulations that including survey weight as a covariate in the propensity score model showed no clear advantage over a survey-weighted propensity score model. Neither approach can consistently outperform the other regarding covariate balance and bias in the treatment effect estimation across various scenarios. Dong et al. [12] evaluated multiple strategies for incorporating survey weights into both propensity score and outcome models, recommending their use at both stages, although they observed that estimates from different strategies did not yield apparent differences in their data application. Lenis et al.[13] found through simulations that incorporating survey weights in the propensity score model does not necessarily improve the estimation of population-level treatment effect using matching, especially when sufficient population-level balance in confounders can already be achieved. They also showed through simulations that incorporating survey weights in propensity score estimation can reduce bias when the missingness in outcome is correlated with survey weights.[13]
@@
Regarding propensity score weighting specifically, Ridgeway et al.[4] provided a theoretical justification for the use of a survey-weighted propensity score model in obtaining a consistent estimator of the population-level average treatment effect. They proved that using survey weights in both stages leads to consistent and robust population average treatment effect estimators under various generating mechanisms. Salerno et al. [14] recently explored a complex survey mechanism in health disparities research, where survey weights depend on the group variable of interest under comparison, such as race. They further revealed that whether the propensity score should be survey-weighted depends on whether the available survey weights are tied to comparison variables such as treatment or exposure. Whereas Ridgeway et al.[4] derived the results and performed simulations for the continuous outcome scenario, Yang et al.[15] extended the results to a binary outcome and evaluated the performance of different ways to incorporate survey weights in the propensity score model and the outcome model. Both Ridgeway et al.[4] and Yang et al.[15] considered the potentially differentiated practice of incorporating survey weights when estimating the average treatment effect (ATE) and the average treatment effect of the treated (ATT).
@@
In general, the emerging message from the existing literature is that the use of survey weights at the propensity score model stage and the outcome model stage can potentially be beneficial in reducing the bias of population-level inference. However, a consensus on the best practice and the underlying justifications for using survey weights to target different causal estimands of interest have not been fully elucidated. This paper aims to contribute to the literature on propensity score weighting in several ways. First, the target estimands in the prior development on survey observational data methods were limited to ATE and ATT, and have not covered other weighted average treatment effect estimands.[9] We address this by supplying a unified framework that incorporates survey weights into the general framework of balancing weights.[9] Second, as a further extension, we detail the use of survey weights for augmented weighting estimators under the balancing weights framework, which can potentially improve efficiency over weighting alone by leveraging information from a posited outcome regression function. Third, we developed closed-form variance estimators that accommodate survey weights in each estimator we considered for computationally efficient inference. Lastly, given that no consensus has been reached on the best use of survey weights in propensity score weighting analysis under various real-world scenarios (such as varying levels of covariate overlap and model misspecification), we conduct extensive simulations to shed light on the operating characteristics of these methods and provide practical recommendations.
@@
The rest of this paper is organized as follows. Section 2 defines the notation and assumptions for the potential outcomes framework in the context of survey observational data. Under the balancing weight framework, this section also presents the weighting estimator and discusses the role of survey weights in addition to balancing weights. In Section 3, we present three augmented estimators incorporating survey weights and further provide the the associated sandwich variance structure for large-sample inference. In Section 4, we conduct simulations to examine the performance of different estimators incorporating propensity score weights for targeting different estimands. Section 5 compares the performance of the different estimation procedures in two real data examples, and Section 6 concludes.
@@ 2.1 Definitions and assumptions
Consider a finite sample of size $n$ selected from a original population of size $N$. Each individual in the sample is indexed by $i$, where $i=1,\dots,n$. We consider a $p$-dimensional vector for pre-treatment covariates for each individual $\mathbf{X}_{i}$, and a binary exposure variable $Z_{i}=z$ that indicates treatment ($z=1$) or control ($z=0$) group affiliation. Under the Stable Unit Treatment Value Assumption (SUTVA), the potential outcomes $Y_{i}(z)$ for $z\in\{0,1\}$ are assumed to be well-defined, and the observed outcome for each individual is $Y_{i}=Z_{i}Y_{i}(1)+(1-Z_{i})Y_{i}(0)$. For estimating population effects using our sample, we define the survey indicator $S_{i}$ where $S_{i}=1$ denotes selection into the sample and $S_{i}=0$ otherwise. Following Salerno et al.,[14] we distinguish two sampling regimes based on when sampling occurs relative to treatment assignment. In a retrospective design, sampling occurs after treatment has been assigned in the population, and the sampling mechanism may depend on both treatment status and covariates. This yields treatment-specific sampling probabilities
$p_{z}(\mathbf{x}_{i})\equiv P(S_{i}=1\mid Z_{i}=z,\mathbf{X}_{i}=\mathbf{x}_{i})~~~\text{for}~~~z\in\{0,1\},$
with the positivity constraint $0<p_{z}(\mathbf{X}_{i})<1$. In a prospective design, sampling occurs before treatment assignment, so the sampling mechanism depends only on covariates with sampling probability
$p(\mathbf{x}_{i})\equiv P(S_{i}=1\mid\mathbf{X}_{i}=\mathbf{x}_{i}),$
which satisfies $0<p(\mathbf{X}_{i})<1$. Under the prospective sampling, $S\perp Z\mid\mathbf{X}$ implies $p_{1}(\mathbf{x})=p_{0}(\mathbf{x})=p(\mathbf{x})$. We further define the population-level propensity score $e_{\text{sp}}(\mathbf{x}_{i})\equiv P(Z_{i}=1\mid\mathbf{X}_{i}=\mathbf{x}_{i})$ and the sample-level propensity score $e_{\text{fp}}(\mathbf{x}_{i})\equiv P(Z_{i}=1\mid\mathbf{X}_{i}=\mathbf{x}_{i},S_{i}=1)$, both satisfying positivity constraints $0<e_{\text{sp}}(\mathbf{X}_{i})<1$ and $0<e_{\text{fp}}(\mathbf{X}_{i})<1$ on the covariate support.
In a retrospective design, we assume conditional independence (or unconfoundedness) for the treatment assignment, $Y(z)\perp Z\mid\mathbf{X}$, and for the post-exposure sampling process, $Y(z)\perp S\mid Z,\mathbf{X}$. These conditions imply that confounding bias arises solely from $\mathbf{X}$, but selection bias can be caused by $Z$ and $\mathbf{X}$. In contrast, a prospective design requires slightly different assumptions. Under noninformative prospective sampling, we assume that potential outcomes are independent of sampling given only covariates, $Y(z)\perp S\mid\mathbf{X}$, and that sampling and treatment are conditionally independent, $S\perp Z\mid\mathbf{X}$. Furthermore, we assume that treatment assignment is unconfounded within the sample such that $Y(z)\perp Z\mid S=1,\mathbf{X}$. These conditions imply that both selection bias and subsequent treatment assignment bias arise solely through $\mathbf{X}$, simplifying the identification strategy.
Under the retrospective designs, the relationship between population and sample propensity scores follows directly from Bayes’ rule:
$e_{\text{fp}}(\mathbf{x})=\frac{p_{1}(\mathbf{x})\,e_{\text{sp}}(\mathbf{x})}{p_{1}(\mathbf{x})\,e_{\text{sp}}(\mathbf{x})+p_{0}(\mathbf{x})\{1-e_{\text{sp}}(\mathbf{x})\}}=\operatorname{expit}\!\left(\operatorname{logit}e_{\text{sp}}(\mathbf{x})+\log\displaystyle\frac{p_{1}(\mathbf{x})}{p_{0}(\mathbf{x})}\right).$ (1)
This identity reveals how treatment-dependent sampling tilts the population propensity scores: the log conditional sampling ratio $\log\{p_{1}(\mathbf{x})/p_{0}(\mathbf{x})\}$ shifts the population-level propensity score to produce the sample-level propensity score. Large discrepancies in sampling rates, either $p_{1}(\mathbf{x})\gg p_{0}(\mathbf{x})$ or $p_{1}(\mathbf{x})\ll p_{0}(\mathbf{x})$, can drive $e_{\text{fp}}(\mathbf{x})$ toward the boundaries, threatening sample-level positivity even when population-level positivity holds. In contrast, under prospective designs, we have $p_{1}(\mathbf{x})=p_{0}(\mathbf{x})=p(\mathbf{x})$, so Equation (1) reduces to $e_{\text{fp}}(\mathbf{x})\equiv e_{\text{sp}}(\mathbf{x})$. This distinction clarifies when survey-weighted propensity scores are necessary for consistent inference: under retrospective (informative) sampling, survey-weighted estimation is required to recover $e_{\text{sp}}(\mathbf{x})$; under prospective (non-informative) sampling where $e_{\text{fp}}(\mathbf{x})=e_{\text{sp}}(\mathbf{x})$, unweighted estimation is theoretically valid.
The above relationship carries implications for certain estimators we will introduce later. To complete the definition of notation, we define $r_{z}(\mathbf{x})=p(\mathbf{x})/p_{z}(\mathbf{x})$ for $z\in\{0,1\}$. From Equation (1) and $p(\mathbf{x})=e_{\text{sp}}(\mathbf{x})p_{1}(\mathbf{x})+\{1-e_{\text{sp}}(\mathbf{x})\}p_{0}(\mathbf{x})$, we arrive at $r_{1}(\mathbf{x})=e_{\text{sp}}(\mathbf{x})/e_{\text{fp}}(\mathbf{x})$ and $r_{0}(\mathbf{x})=\{1-e_{\text{sp}}(\mathbf{x})\}/\{1-e_{\text{fp}}(\mathbf{x})\}$, which together yield $p_{1}(\mathbf{x})/p_{0}(\mathbf{x})=r_{0}(\mathbf{x})/r_{1}(\mathbf{x})$. Here $r_{1}(\mathbf{x})>1$ indicates undersampling of treated units relative to their population prevalence, while $r_{1}(\mathbf{x})<1$ indicates oversampling. The relationship $p(\mathbf{x})=r_{z}(\mathbf{x})p_{z}(\mathbf{x})$ provides a bridge for estimating marginal sampling probabilities when only treatment-specific ones are available from survey data, under retrospective designs. As Salerno et al.[14] highlighted, the distinction between prospective and retrospective studies corresponds to two different factorizations of $P(S=1,Z=z\mid\mathbf{X})$: one conditional on covariates alone versus one conditional on both treatment and covariates. In this paper, we primarily focus on the retrospective design, but provide the counterparts of our methods under prospective design in Web Appendix A.6.
We define the conditional average treatment effect as $\tau(\mathbf{x})=E[Y(1)-Y(0)\mid\mathbf{X}=\mathbf{x}]$, which is the building block for marginal estimands depending on the covariate distribution. Although prior work has primarily focused on the population average treatment effect (PATE) and the population average treatment effect of the treated (PATT),[4, 15, 14] we follow the framework of balancing weights [9] that characterizes a general class of weighted distributions of covariates with an extension to survey observational studies. Specifically, the density of the target population is represented as $g(\mathbf{X})=f(\mathbf{X})h(\mathbf{X})$, where $f(\mathbf{X})$ signifies the marginal density of covariates of the original external population, and $h(\mathbf{X})$, known as the tilting function, characterizes the how that original external population is tilted to generate the new target population. The general class of population weighted average treatment effect (PWATE) is given by
@@
In this equation, the expectations in both the numerator and the denominator are taken over the target population. The tilting function $h(\mathbf{X})$ is a non-negative function of covariates and satisfies $E\left[h(\mathbf{X})\right]>0$. When $h(\mathbf{X})=1$ and $h(\mathbf{X})\propto e_{\text{sp}}(\mathbf{X})$, the PWATE estimand $\tau_{h}$ reduces to PATE and PATT, respectively.
2.2 Incorporating survey weights into propensity score weighting with balancing weights
Under the balancing weights framework, the tilting function $h(\mathbf{X})=g(\mathbf{X})/f(\mathbf{X})$ is used to tilt the covariate distribution in the population $f(\mathbf{X})$ to that in the target population $g(\mathbf{X})$. Since our goal is to derive the expressions of various estimands for target populations in the original population, we first provide a general identification formula of this class of estimands $\tau_{h}(\mathbf{\mathbf{X}})$ considering the average treatment effect over the target population in the original population. Under the retrospective design, Equation (2) can be expressed as,
$\tau_{h}$ $=\frac{E\left[\displaystyle\frac{h(\mathbf{X})ZY}{p_{1}(\mathbf{X})e_{\text{sp}}(\mathbf{X})}\Bigg|S=1\right]}{E\left[\displaystyle\frac{h(\mathbf{X})Z}{p_{1}(\mathbf{X})e_{\text{sp}}(\mathbf{X})}\Bigg|S=1\right]}-\frac{E\left[\displaystyle\frac{h(\mathbf{X})(1-Z)Y}{p_{0}(\mathbf{X})\left(1-e_{\text{sp}}(\mathbf{X})\right)}\Bigg|S=1\right]}{E\left[\displaystyle\frac{h(\mathbf{X})(1-Z)}{p_{0}(\mathbf{X})\left(1-e_{\text{sp}}(\mathbf{X})\right)}\Bigg|S=1\right]}.$ (3)
When the expectations are replaced with sample means, an unbiased estimator of $\tau_{h}$ in the form of a weighted difference is expressed as